OpenStudy (anonymous):

Eli owes $605. The first week he will pay back $20. The second week he will pay back $29. The third week he will pay back $38. If Eli continues to pay the money back using the same sequence, how many weeks will it take him to pay back the entire amount?

OpenStudy (anonymous):

sequence is just 20+9x, where x is weeks so 605=20+9x Solve for x x=65